OpenStudy (anonymous):

@Mathematics x f(x) -6 8 -5 0 -4 -6 -3 -10 -2 -12 -1 -12 0 -10 1 -6 2 0 Thomas creates the table shown to represent the function ƒ(x) = x2 + 3x - 10. Determine the zeros of the equation. A) x = -5 and x = 2 B) x = 0 and x = -3 C) x = 0 and x = -10 D) x = -2 and x = -1

OpenStudy (anonymous):

\[x^2+3x-10=0\] \[(x+5)(x-2)=0\] \[x+5=0\iff x=-5\] \[x-2=0\iff x=2\]

OpenStudy (anonymous):

but since the table actually already includes 2 0 -5 0 you had the answer already given

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

that would be A